In a family of 11 children, what is the probability that there will be more boys than girls?






1
Expert's answer
2022-01-25T06:07:43-0500

"n=11"

To get more boys than girls means at least 6 boys should be there for this condition


"p=P(Boy) = 0.5 \\"


"q=1-p=1-P(boy)=P(Girl) = 0.5"

Probability for at least 6 boys:


"P(X\u22656) = P(X=6) + P(X=7) + P(X+8)"


"+ P(X=9) + P(X=10) + P(X=11)"

"=\dbinom{11}{6}(0.5)^6(0.5)^{11-6}+\dbinom{11}{7}(0.5)^7(0.5)^{11-7}"

"+\dbinom{11}{8}(0.5)^8(0.5)^{11-8}+\dbinom{11}{9}(0.5)^9(0.5)^{11-9}"

"+\dbinom{11}{10}(0.5)^{10}(0.5)^{11-10}+\dbinom{11}{11}(0.5)^{11}(0.5)^{11-11}"

"=(462+330+165+55+11+1)(0.5)^{11}"

"=1024(0.5)^{11}=0.5"

There is 50% probability that there is more boys than girls.